OpenStudy (anonymous):

write the trig expression in terms of sine and cosine, then simplify: Sin(x)Sec(x)

OpenStudy (solomonzelman):

\[we~~~know~~~that~~~~~~\sf Sec(x)=\frac{1}{Cos(x)}\] so your equation turns into \[\sf Sin(x)\times \frac{1}{Cos(x)}\] and that's the same as \[\sf \frac{Sin(x)}{Cos(x)}\] can you tell me the answer now, or you need more help? (don't hesitate to reply!)

OpenStudy (anonymous):

Tan(x) Thank you, I don't understand these trig functions at all!

OpenStudy (solomonzelman):

What exactly do you not understand about them? You mean this type of problems that you don't get? or how to find a tan of an angle or what?

OpenStudy (anonymous):

These types of problems.

OpenStudy (solomonzelman):

Here are some identities. Knowing these identites and applying them will help you. the more you practice the better you will get. (I also went through that) \(\LARGE\color{blue}{ \sf Sin(x) }\)\(\Huge\color{blue}{ \sf =\frac{1}{Csc(x)} }\) \(\LARGE\color{blue}{ \sf Csc(x) }\)\(\Huge\color{blue}{ \sf =\frac{1}{Sin(x)} }\) \(\LARGE\color{blue }{ \sf Cos(x) }\)\(\Huge\color{blue}{ \sf =\frac{1}{Sec(x)} }\) \(\LARGE\color{blue}{ \sf Sec(x) }\)\(\Huge\color{blue}{ \sf =\frac{1}{Cos(x)} }\) \(\LARGE\color{blue}{ \sf Tan(x) }\)\(\Huge\color{blue}{ \sf =\frac{1}{Cot(x)} }\) \(\LARGE\color{blue}{ \sf Cot(x) }\)\(\Huge\color{blue}{ \sf =\frac{1}{Tan(x)} }\) also te Pythagorean identities. HERE....

OpenStudy (solomonzelman):

oops, I posted the blue once awkwardly.

OpenStudy (anonymous):

Its okay! So how about the problem tan^2(x)-sec^2(x)... I have it to Sin^2(x)/Cos^2(x)-1/cos^2(x)... but I have no clue where to go from there...:(

OpenStudy (solomonzelman):

combine the fractions Sin^2(x)/Cos^2(x)-1/cos^2(x) [ Sin^2x-1 ] / Cos^2x [-Cos^2x / Cos^2x] -1